Jeremy already drove 40 miles from home and continued at an average of 60 mph. Jared left the house and drove an average of 70 mph. How many hours will it take for Jeremy and Jared to meet up? Write a system of equations and solve!
Thank you!

Answer by Alan3354(69443)

Jeremy already drove 40 miles from home and continued at an average of 60 mph. Jared left the house and drove an average of 70 mph. How many hours will it take for Jeremy and Jared to meet up?
----------
Jeremy is 40 miles ahead.
Jared "gains on him" at 10 mi/hr (70-60)
----
d = r*t
t = d/r = 40/10
t = 4 hours